### v1

@@ Abstract
Autonomous agents that execute human tasks by controlling computers can enhance human productivity and application accessibility. Yet, progress in this field will be driven by realistic and reproducible benchmarks. We present AndroidWorld, a fully-functioning Android environment that provides reward signals for 116 programmatic task workflows across 20 real-world Android apps. Unlike existing interactive environments, which provide a static test set, AndroidWorld dynamically constructs tasks that are parameterized and expressed in natural language in unlimited ways, thus enabling testing on a much larger and realistic suite of tasks. Reward signals are derived from the computer’s system state, making them durable across task variations and extensible across different apps.
To demonstrate AndroidWorld’s benefits and mode of operation, we introduce a new computer control agent, M3A. M3A can complete 30.6% of the AndroidWorld’s tasks, leaving ample room for future work. Furthermore, we adapt a popular desktop web agent to work on Android, which we find to be less effective on mobile, suggesting future research is needed to achieve universal, cross-domain agents. Finally, we conduct a robustness analysis by testing M3A against a range of task variations on a representative subset of tasks, demonstrating that variations in task parameters can significantly alter the complexity of a task and therefore an agent’s performance, highlighting the importance of testing agents under diverse conditions. AndroidWorld and the experiments in this paper are available at https://github.com/google-research/android_world.
@@ 1 Introduction
Autonomous agents that interpret human natural language instructions and operate computing devices can provide enormous value by automating repetitive tasks, augmenting human intelligence, and accomplishing complex workflows. The enthusiasm for building such agents is evident by the growing number of computer control agent systems [41, 9, 70, 24, 22, 16] and code repositories [13, 1, 2, 57]. Yet, most existing approaches evaluate agents with proxy metrics, by comparing an agent’s trajectory to a previously collected human demonstration [28, 5, 9, 41, 63, 68, 33, 67, 61]. This kind of measurement is not representative of real-world performance [41, 70] because it does not reflect that there are usually multiple correct paths to solve a task, environments may behave non-deterministically (e.g., the environment freezes), and agents can dynamically learn from mistakes to correct their actions [46, 31, 27, 39]. Towards more accurate evaluation, emerging testing environments [73, 24, 64, 59] place agents in interactive environments and measure their task completion rates across a suite of tasks.
Reward signals are quantitative metrics that indicate functional correctness of a task, i.e. is the stated goal achieved? For example, for the task “Send a text message to Jane confirming I’ll be there”, a positive reward indicates the relevant message has been sent. Yet, real-world environments (apps and websites), unlike simulations [49, 47] and games [38, 48, 55], do not naturally provide explicit reward signals. The internals of most apps and websites are hidden from developers making it difficult to program rewards. Additionally, such systems are seemingly infinitely expansive, spanning millions of continuously-changing apps and websites. All these factors make it hard to establish reliable and durable reward functions.
One way to obtain reliable reward signals is to leverage human judgment [41, 70, 39, 23]. This allows one to conduct one-off evaluations on a diverse suite of tasks, however each evaluation run requires significant time and costly manual review. LLM-based “judges” used to evaluate natural language generation on open-ended questions [7, 71, 32] are emerging as a solution to dynamic real-world environments [10, 35] and while showing promise also for computer control agents [39, 16] they remain imperfect evaluators. As a result, existing testing environments for computer control agents that provide automated reward signals are limited in their real-world diversity and scale. They cover mostly single-app tasks on a small set of domains (1–6 websites [64, 73, 24]) and evaluate agents on static test sets [59, 37], instead of across across varied conditions and inputs which are found in real-world scenarios.
@@
To tackle these challenges, we introduce and release AndroidWorld, a robust agent environment for Android that offers realism and scalability for developing and evaluating computer control agents. AndroidWorld is an open environment consisting of a fully-functioning Android OS, allowing access to millions of Android apps and the Web. Crucially, it provides a highly reproducible task suite for developing and evaluating computer control agents, consisting of 116 tasks across 20 apps. Similarly to the popular MiniWoB++ [45] benchmark, each task is dynamically instantiated using randomly-generated parameters, challenging agents with millions of unique task goals and conditions. While MiniWob++ consists of simple, synthetic websites, AndroidWorld supports real-world Android applications. To make reward signals durable using real applications, AndroidWorld’s key insight is to leverage the extensive and consistent state management capabilities of the Android operating system, using the same mechanisms that the apps themselves utilize. The dynamic nature of AndroidWorld’s tasks is not only a test of adaptability for evaluation purposes, but it can also spawn new research on online learning algorithms for computer control agents.
@@
To demonstrate AndroidWorld’s usefulness as a benchmark, we build and release a new multi-modal agent, M3A (Multimodal Autonomous Agent for Android), and establish state-of-the-art results on AndroidWorld. We analyze M3A’s performance using both multimodal and text-only input, and we show that multimodal perception significantly improves agent performance. On AndroidWorld, M3A achieves a 30.6% success rate. This performance surpasses that of a web agent adapted for Android but remains significantly lower than the human success rate of 80.0%. In pursuit of building robust computer control agents, our study extends to a series of real-world condition tests, where we show performance varies significantly across variations in intent parameters, phrasing, and environmental configurations.
@@ 2.1 Interactive evaluation environments
Effective evaluation of autonomous agents requires environments that not only mimic real-world scenarios but also provide immediate reward signals upon successful task completion [41, 9, 3, 42, 6]. MiniWoB++ [45, 31] offers a lightweight framework of small, synthetic HTML pages with parameterized tasks which allow for unlimited task variability. WebShop [64] provides a simulated e-commerce environment, whereas WebArena [73] and VisualWebArena [24] consist of simulated websites across up to four domains. OSWorld [59] provides an interface and programmatic rewards across nine apps for desktop OSes. GAIA [37] is a static dataset that tests an agent’s ability to interact with live web environments. MMInA [69] is a multihop and multimodal benchmark designed to evaluate agents for compositional Internet tasks. B-MoCA [25] is the only interactive testing environment that exists today for mobile. It supports 31 tasks whose success is determined by comparing an agent’s execution with logs from human demonstrations using simple regex expressions. Unlike MiniWoB++, all these recent frameworks use a fixed set of evaluation tasks where initial conditions and goals are statically defined, limiting adaptability. As has been observed in other domains such as reinforcement learning [17, 40, 8], the current output of foundation models is stochastic and highly sensitive to their inputs [20, 34, 30, 43], motivating the need to test them under varying conditions.
@@
Additionally, other studies leverage human evaluation [41, 70, 4] for tasks where automatic evaluation is not available. Lastly, emerging research [39, 16] explores the potential of multimodal models to generalize agent evaluations to new settings, though this area requires further research to achieve accuracy comparable to manually coded rewards.
@@ 2.3 Interactive agents
Prior to today’s foundation models, traditional approaches to developing user interface-operating agents primarily used reinforcement learning and behavioral cloning to simulate interactions like mouse clicks and keyboard typing [28, 31, 14, 19]. More recent work tends to leverage off-the-shelf foundational models [50, 51, 52] with in-context learning (ICL) and fine-tuning applied to mobile [41, 18, 56, 61, 68, 4], desktop web [70, 9, 73, 24], and desktop OS [58, 66, 59]. Recent work explores agents that reflect on system state [46, 65, 36] by leveraging exploration, self-evaluation, and retry-capabilities enabling continual learning and adaptation [27, 63, 39, 58].
@@ 3.1 Android for autonomous agents
Android is an ideal environment for developing autonomous agents. It is the most widely used operating system globally [^2] https://gs.statcounter.com/os-market-share. and is highly flexible for research, while providing an open world of the Web [^3] Mobile is the most popular platform for accessing the web; https://gs.statcounter.com/platform-market-share/desktop-mobile/worldwide/ and over 2M applications for agents to operate in. Using emulation, the Android environment is easy to deploy, does not require specialized hardware, and can be run on a laptop. Android Virtual Devices (AVDs), or emulator images, are well suited for research because they are self-contained, easily distributable, and configurable.
@@ 3.2 The observation and action space
AndroidWorld provides an interface for agents to receive observations and execute actions on Android. It uses AndroidEnv [53] and the Android Device Bridge to facilitate interaction between Android and the agent. The observation space consists of a full-resolution screenshot and a UI tree representation developed for accessibility purposes. The action space is similar to that which humans use, consisting of gestures (i.e., tapping, long-press, and swiping), typing, and navigation buttons (i.e., go home and go back). In addition to these naturalistic actions, AndroidWorld exposes a limited set of function calling APIs, such as send_text_message, to help agents accomplish goals. Appendix A provides more details on the observation format and action space.
@@ 3.3 Reproducible and parameterized tasks
Figure 2: Human raters assessed task difficulty (2(a)), duration (2(b)), and category (2(c)). Each rater performed the tasks assigned to them, then assigned a difficulty level and selected relevant tags from a predefined list. They also estimated the number of steps required to complete each task, using the action space available to an agent.
AndroidWorld consists of a suite of 116 tasks, spread across 20 diverse applications (see Appendix B for more details). These tasks simulate practical, everyday activities, including note-taking, scheduling appointments, communicating through messaging, and interacting with system utilities. The suite consists of open-source apps and built-in Android system apps, such as Settings and Contacts. As rated by humans, the tasks vary in difficulty, duration, and categories (Figure 2).
To achieve a high degree of reproducibility in real-world scenarios, AndroidWorld precisely controls the OS and app state in several ways. The Android OS is fixed, consisting of a Pixel 6 emulator running Android 13 with a fixed time on the date October 15th, 2023. All applications in AndroidWorld are fully-functional, open-source apps, besides the OS-level apps included with Android. For the open-source apps, AndroidWorld maintains a constant environment by installing a fixed version of each app, acquired from F-Droid.  [^4] https://f-droid.org/OS-level apps’ versions are determined by the Android OS, which is also fixed. AndroidWorld only utilizes apps that do not require login/authentication, can function offline, and store their application data on device.
In addition to managing the states of apps and operating systems, AndroidWorld precisely defines and controls the state during task execution. Each task has its own unique setup, reward determination logic, and teardown procedures (see Appendix B.2 for a detailed example), ensuring a fully reproducible suite of tasks. This careful state management not only ensures reproducibility but also allows for easy customization of tasks through parametrization. Task parameters, initialized randomly at the start of each task based on a controlled random seed, dictate the initial state and influence reward outcomes. Similarly to MiniWoB++ [45, 29], AndroidWorld consists of a practically infinite set of varying initial conditions and success criteria. This approach provides more granular analyses of agents’ adaptability — a vital attribute for real-world deployment. Beyond testing agent robustness, the dynamic construction of tasks supports the use of online learning methodologies, particularly reinforcement learning [45, 29, 19, 14]. It also simplifies the generation of distinct train/test datasets, facilitating supervised learning experiments [19, 44, 12].
@@
Create a new note in Markor named {file_name} with the following text: {text}. Share the entire content of the note with the phone number {number} via SMS. (note_created(file_name, text) + message_exists(phone_number, message)) / 2.0
@@ 3.4 Durable rewards from system state
AndroidWorld provides reward signals by managing application state using the Android Debug Bridge (adb). With the adb tool AndroidWorld has complete access to system resources including the file system, application databases, and system settings. Determining reward signals from system state has several benefits. It is highly accurate because an application’s state can be quickly inspected and manipulated using the same mechanisms that the app itself utilizes. Using the underlying system state is much more durable than matching superficial UI changes. Additionally, it facilitates easy re-use across disparate apps, which tend to use the same underlying caching mechanisms. For instance, logic for checking existence of a specific file is used across many unrelated applications, including those for file management, note-taking, and media playback. For applications leveraging SQLite databases, a common pattern, AndroidWorld implements evaluators that verify the existence of new and deleted rows. Table 2 shows examples of the validators in AndroidWorld. For more examples see Table 5 in the Appendix.
@@ 3.6 Integrating MiniWob++
To demonstrate AndroidWorld’s extensibility, we implement the MiniWoB++ in the AndroidWorld framework and term it MobileMiniWoB++. Each MiniWoB++ task is instantiated using the standard AndroidWorld interface, inheriting from TaskEval base class, and contains methods like initialize_state and is_successful. Since MiniWoB++ leverages JavaScript for task configuration and success detection, we built a WebView app to communicate between Python and the app. For instance, the is_successful method of each task retrieves the reward value from the WebView app via an Android intent.
MobileMiniWoB++ introduces modifications in both observations and actions compared to the original benchmark. For example, HTML5 <input> elements are natively rendered with native Android UI widgets like the date-picker (see Figure  4 in the Appendix), enhancing the realism of the tasks. MobileMiniWoB++ uses the same observation space as the Android tasks (accessibility tree and screenshot). Notably, it does not include the DOM as in the original implementation. The action space from AndroidWorld is retained. We manually review and test each task to ensure they are solvable. We excluded twelve of the original tasks that failed to render correctly on Android, presented compatibility issues with the touch interface, or required near real-time interaction, which poses challenges on emulators. Overall, AndroidWorld supports 92 MiniWoB++ tasks.

### v2

@@ Abstract
Autonomous agents that execute human tasks by controlling computers can enhance human productivity and application accessibility. However, progress in this field will be driven by realistic and reproducible benchmarks. We present AndroidWorld, a fully functional Android environment that provides reward signals for 116 programmatic tasks across 20 real-world Android apps. Unlike existing interactive environments, which provide a static test set, AndroidWorld dynamically constructs tasks that are parameterized and expressed in natural language in unlimited ways, thus enabling testing on a much larger and more realistic suite of tasks. Reward signals are derived from the computer’s system state, making them durable across task variations and extensible across different apps.
To demonstrate AndroidWorld’s benefits and mode of operation, we introduce a new computer control agent, M3A. M3A can complete 30.6% of AndroidWorld’s tasks, leaving ample room for future work. Furthermore, we adapt a popular desktop web agent to work on Android, which we find to be less effective on mobile, suggesting future research is needed to achieve universal, cross-domain agents. Finally, we conduct a robustness analysis by testing M3A against a range of task variations on a representative subset of tasks, demonstrating that variations in task parameters can significantly alter a task’s complexity and, consequently, an agent’s performance, highlighting the importance of testing agents under diverse conditions. AndroidWorld and the experiments in this paper are available at https://github.com/google-research/android_world.
@@ 1 Introduction
Autonomous agents that interpret human natural language instructions and operate computing devices can provide enormous value by automating repetitive tasks, augmenting human intelligence, and accomplishing complex workflows. The enthusiasm for building such agents is evident by the growing number of computer control agent systems [45, 10, 78, 27, 25, 19] and code repositories [16, 1, 2, 63, 66]. Yet, most existing approaches evaluate agents with proxy metrics, by comparing an agent’s trajectory to a previously collected human demonstration [32, 5, 10, 45, 70, 76, 37, 74, 68]. This kind of measurement is not representative of real-world performance [45, 78] because it does not reflect that there are usually multiple correct paths to solve a task, environments may behave non-deterministically (e.g., the environment freezes), and agents can dynamically learn from mistakes to correct their actions [50, 35, 31, 43]. Towards more accurate evaluation, emerging testing environments [71, 81, 27, 11, 29, 65] place agents in interactive environments and measure their task completion rates across a suite of tasks.
Reward signals are quantitative metrics that indicate functional correctness of a task, i.e. is the stated goal achieved? For example, for the task “Send a text message to Jane confirming I’ll be there”, a positive reward indicates the relevant message has been sent. Yet, real-world environments (apps and websites), unlike simulations [54, 51] and games [42, 52, 60, 62, 53], do not naturally provide explicit reward signals. The internals of most apps and websites are hidden from developers making it difficult to program rewards. Additionally, such systems are seemingly infinitely expansive, spanning millions of continuously-changing apps and websites. All these factors make it hard to establish reliable and durable reward functions.
One way to obtain reliable reward signals is to leverage human judgment [45, 78, 43, 26] to conduct one-off evaluations on a diverse suite of tasks; however, each evaluation run requires both significant time and costly manual review. LLM-based “judges” used to evaluate natural language generation models  [8, 79, 36] are emerging as a solution for evaluating agents in dynamic real-world environments [12, 39], and while showing promise also for computer control agents [43, 19] they remain imperfect evaluators. As a result, existing testing environments for computer control agents that provide automated reward signals are limited in their real-world diversity and scale. They cover mostly single-app tasks on a small set of domains (1–6 websites [71, 81, 27]) and evaluate agents on static test sets [65, 41], instead of across varied conditions and inputs which are found in real-world scenarios.
@@
To tackle these challenges, we introduce and release AndroidWorld, a robust agent environment for Android that offers realism and scalability for developing and evaluating computer control agents. AndroidWorld is an open environment consisting of a fully functional Android OS, allowing access to millions of Android apps and the Web. Crucially, it provides a highly reproducible task suite for developing and evaluating computer control agents, consisting of 116 tasks across 20 apps. Similarly to the popular MiniWoB++ [49] benchmark, each task is dynamically instantiated using randomly-generated parameters, challenging agents with millions of unique task goals and conditions. While MiniWob++ consists of simple, synthetic websites, AndroidWorld leverages real-world Android applications. To ensure that reward signals are durable when using real applications, AndroidWorld’s key insight is to leverage the extensive and consistent state management capabilities of the Android operating system, using the same mechanisms that the apps themselves utilize. The dynamic nature of AndroidWorld’s tasks is not only a test of adaptability for evaluation purposes, but it can also spawn new research on online learning algorithms for computer control agents.
@@
To demonstrate AndroidWorld’s usefulness as a benchmark, we build and release a new multi-modal agent, M3A (Multimodal Autonomous Agent for Android), and establish state-of-the-art results on AndroidWorld. We analyze M3A’s performance using both multimodal and text-only input, and we observe that while multimodal perception can improve performance in some cases, it generally does not outperform the text-only approach. On AndroidWorld, M3A achieves a 30.6% success rate, which surpasses that of a web agent adapted for Android but remains significantly lower than the human success rate of 80.0%. In pursuit of building robust computer control agents, our study includes comprehensive tests under varied real-world conditions, demonstrating significant performance variations primarily driven by changes in intent parameters.
@@ 2.1 Interactive evaluation environments
Effective evaluation of autonomous agents requires environments that not only mimic real-world scenarios but also provide immediate reward signals upon successful task completion [45, 10, 3, 46, 6]. MiniWoB++ [49, 35] offers a lightweight framework of small, synthetic HTML pages with parameterized tasks which allow for unlimited task variability. WebShop [71] provides a simulated e-commerce environment, whereas WebArena [81] and VisualWebArena [27] consist of simulated websites across up to four domains. WorkArena [11] consists of 29 tasks for enterprise software. OSWorld [65] provides an interface and programmatic rewards across nine apps for desktop OSes. GAIA [41] is a static dataset that tests an agent’s ability to interact with live web environments. MMInA [77] is a multihop and multimodal benchmark designed to evaluate agents for compositional Internet tasks. B-MoCA [29] is the only interactive testing environment that exists today for mobile. It supports 31 tasks whose success is determined by comparing an agent’s execution with logs from human demonstrations using simple regex expressions. Unlike MiniWoB++, these frameworks use a fixed set of evaluation tasks with static initial conditions and goals, limiting adaptability. As has been observed in other domains such as reinforcement learning [20, 44, 9], the current output of foundation models is stochastic and highly sensitive to their inputs [23, 38, 34, 47], motivating the need to test them under varying conditions.
@@
Other studies leverage human evaluation [45, 78, 4] for tasks where automatic evaluation is not available. Lastly, emerging research [43, 19] explores the potential of multimodal models to generalize agent evaluations to new settings, though this area requires further research to achieve accuracy comparable to manually coded rewards.
@@ 2.3 Interactive agents
Prior to today’s foundation models, traditional approaches to developing user interface-operating agents primarily used reinforcement learning and behavioral cloning to simulate interactions like mouse clicks and keyboard typing [32, 35, 17, 22]. More recent work tends to leverage off-the-shelf foundational models [55, 56, 57] with in-context learning (ICL) and fine-tuning applied to mobile [45, 21, 61, 68, 76, 4, 75], desktop web [78, 10, 81, 27, 7, 28], and desktop OS [64, 73, 65]. Recent work explores agents that reflect on system state [50, 72, 40] by leveraging exploration, self-evaluation, and retry-capabilities enabling continual learning and adaptation [31, 70, 43, 64, 15].
@@ 3.1 Android for autonomous agents
Android is an ideal environment for developing autonomous agents. It is the most widely-used OS globally [^2] https://gs.statcounter.com/os-market-share and is highly flexible for research, while providing an open world of the Web [^3] Mobile is the most popular platform for accessing the web; https://gs.statcounter.com/platform-market-share/desktop-mobile/worldwide/ and over 2M apps for agents to operate in. Using emulation, an Android environment is easy to deploy, does not require specialized hardware, and can be run on a laptop. Android Virtual Devices or emulator images are well suited for research as they are self-contained, easy to distribute, and configurable.
@@ 3.2 The observation and action space
AndroidWorld provides an interface for agents to receive observations and execute actions on Android. It uses AndroidEnv [58] and the Android Device Bridge to facilitate interaction between Android and the agent. The observation space consists of a full-resolution screenshot and a UI tree representation developed for accessibility purposes. The action space is similar to that which humans use, consisting of gestures (i.e., tapping, long-press, and swiping), typing, and navigation buttons (i.e., go home and go back). In addition to these naturalistic actions, AndroidWorld exposes a limited set of function calling APIs, such as send_text_message, to help agents accomplish goals. App. B provides more details on the observation format and action space.
@@ 3.3 Reproducible and parameterized tasks
Figure 2: Annotators performed the tasks assigned to them, assigned a difficulty level (2(a)) and selected relevant category tags from a predefined list (2(b)). They also estimated the number of steps required to complete each task (2(b)), using the action space available to an agent.
AndroidWorld consists of a suite of 116 tasks, spread across 20 diverse applications (see App. C for more details). These tasks simulate practical, everyday activities, including note-taking, scheduling appointments, communicating through messaging, and interacting with system utilities. The suite consists of open-source apps and built-in Android system apps, such as Settings and Contacts. As rated by humans, the tasks vary in difficulty, duration, and categories (Figure 2).
To achieve a high degree of reproducibility in real-world scenarios, AndroidWorld precisely controls the OS and app states in several ways. The Android OS is fixed, consisting of a Pixel 6 emulator running Android 13 with a fixed time on the date October 15th, 2023. All applications in AndroidWorld are fully-functional, open-source apps, besides the OS-level apps included with Android. For the open-source apps, AndroidWorld maintains a constant environment by installing a fixed version of each app, acquired from F-Droid.  [^4] https://f-droid.org/. OS-level apps’ versions are determined by the Android OS, which is also fixed. AndroidWorld only utilizes apps that do not require login/authentication, can function offline, and store their application data on device.
In addition to managing the states of apps and operating systems, AndroidWorld precisely defines and controls the state during task execution. Each task has its own unique setup, reward determination logic, and teardown procedures (see App. C.2 and C.3 for more details), ensuring a fully reproducible suite of tasks. This careful state management not only ensures reproducibility but also allows for easy customization of tasks through parametrization. Task parameters, initialized randomly at the start of each task based on a controlled random seed, dictate the initial state and influence reward outcomes. Similarly to MiniWoB++ [49, 33], AndroidWorld consists of a practically infinite set of varying initial conditions and success criteria. This approach provides more granular analyses of agents’ adaptability — a vital attribute for real-world deployment. Beyond testing agent robustness, the dynamic construction of tasks supports the use of online learning methodologies, particularly reinforcement learning [49, 33, 22, 17]. It also simplifies the generation of distinct train/test datasets, facilitating supervised learning experiments [22, 48, 14].
@@
Create a new note in Markor named {file_name} with the following text: {text}. Share the entire content of the note with the phone number {number} via SMS. (file_exists(file_name, content=text) + message_exists(phone_number, message)) / 2.0
@@ 3.4 Durable rewards from system state
AndroidWorld provides reward signals by managing application state using the Android Debug Bridge (adb). With the adb tool AndroidWorld has complete access to system resources including the file system, application databases, and system settings. Determining reward signals from system state has several benefits. It is highly accurate because an application’s state can be quickly inspected and manipulated using the same mechanisms that the app itself utilizes. Using the underlying system state is much more durable than matching superficial UI changes. Additionally, it facilitates easy re-use across disparate apps, which tend to use the same underlying caching mechanisms. For instance, logic for checking existence of a specific file is used across many unrelated applications, including those for file management, note-taking, and media playback. For applications leveraging SQLite databases, a common pattern, AndroidWorld implements evaluators that verify the existence of new and deleted rows. Table 2 shows examples of the validators in AndroidWorld. For more examples see Table 5.
@@ 3.6 Integrating MiniWob++
To demonstrate AndroidWorld’s extensibility, we implement MiniWoB++ in the AndroidWorld framework and term it MobileMiniWoB++. Each MobileMiniWoB++ task is instantiated using the standard AndroidWorld interface, inheriting from TaskEval base class, and contains methods like initialize_state and is_successful. Since MiniWoB++ leverages JavaScript for task configuration and success detection, we built a WebView app to communicate between Python and the app. For instance, the is_successful method of each task retrieves the reward value from the WebView app via an Android intent.
MobileMiniWoB++ introduces modifications in both observations and actions compared to the original benchmark. For example, HTML5 <input> elements are rendered with native Android UI widgets like the date-picker (see Figure  4), enhancing the realism of the tasks. MobileMiniWoB++ uses the same observation space as the Android tasks (accessibility tree and screenshot). Notably, it does not include the DOM as in the original implementation. The action space from AndroidWorld is retained. We manually review and test each task to ensure they are solvable. We excluded twelve of the original tasks that failed to render correctly on Android, presented compatibility issues with the touch interface, or required near real-time interaction, which poses challenges on emulators. Overall, AndroidWorld supports 92 MiniWoB++ tasks. See App. B.3 for more details.